OpenStudy (anonymous):

According to the general equation for conditional probability, if p (a^b')=1/6 and p(b')=7/12 , what is p(a[b) ?

OpenStudy (kropot72):

P(B) = 1 - P(B') = 1 - (7/12) = 5/12 \[\large P(A \cap B)=\frac{P(A \cap B')}{P(B')} \times \frac{P(B)}{1}\] @Emjay9295 Can you follow so far?

OpenStudy (kropot72):

Plugging values into the last equation we get: \[\large P(A \cap B)=\frac{1\times12\times5}{6\times7\times12}=\frac{5}{42}\] Now we can make use of the following formula \[\large P(A|B)=\frac{P(A \cap B)}{P(B)}\] by plugging in the values that we have found.

OpenStudy (anonymous):

yeah i got what you did so far, but how would i use the 5/42? in the formula?

OpenStudy (anonymous):

Would it be the top part of the formula?

OpenStudy (anonymous):

and the bottom would be 7/12?

OpenStudy (kropot72):

That is correct. 5/42 is the numerator and the denominator is 5/12.

OpenStudy (anonymous):

so from there how would i go about solving that? Would i be looking to divide those two fractions as is? Or would i switch it to a decimal and divide them?

OpenStudy (kropot72):

The bottom (denominator) is P(B) which equals 5/12. \[\large P(A|B)=\frac{5\times12}{42\times5}=you\ can\ calculate\]

OpenStudy (anonymous):

Is 60/210

OpenStudy (kropot72):

Do you have a choice of answers? If so, are they fractions or are they decimal numbers?

OpenStudy (anonymous):

They are fractions A. 3/7 B .2/7 C.4/7 D.5/7

OpenStudy (anonymous):

Which i think it is 2/7 at this point

OpenStudy (kropot72):

Yes, the fraction that you calculated simplifies to 2/7.

OpenStudy (anonymous):

Thank you for your help, it's really appreciated!

OpenStudy (kropot72):

You're welcome :)